OpenStudy (brightemmerald):

What are the vertex and x-intercepts of the graph of the function given below? y = x^2 - 2x - 35 A. Vertex: (0, 0); intercepts: x = -4, -5 B. Vertex: (1, -36); intercepts: x = 7, -5 C. Vertex: (1, -25); intercepts: x = 6, -4 D. Vertex: (7, 5); intercepts: x = 7, 8

OpenStudy (brightemmerald):

@5ever

OpenStudy (mathmale):

From y = x^2 - 2x - 35 you can obtain the y-intercept right away: Just let x=0. What is y? Finding the x-intercepts is just a bit harder: Set y = x^2 - 2x - 35 equal to 0 and find the two roots / zeroes. These represent the x-coordinates of the x-intercepts. Given y = x^2 - 2x - 35, find the vertex using "completing the square." Rewrite y = x^2 - 2x - 35 as y = x^2 - 2x - 35. Focus on the coefficient (-2) of x. What's your next step? If you've done "completing the square" before, please review that.